OpenStudy (anonymous):

If f(x) = x + 4, find f(-7). 3 11 -3 -11

OpenStudy (anonymous):

-3 Just plug in -7 for x and solve and you get -3

OpenStudy (anonymous):

for future reference if f(x)= x+4 if we're solving for f(0) then f(0) = 0+4 if we're solving for f(5) then f(5) = 5+4 if we're solving for f(-10) then f(-10)= -10 +4 see the pattern?